Find an orthogonal matrix P such that PPAP diagonalizes A = - Lozi 13 2 V2 13
Answer #1

13 52 13 (A - 24 = [ 13 2 > We are given- A = [ur Characteristio eqh of A = 1 A-2I(X=0 13-7 h = (2-13) ?_2=0 13-a 22 +169 -26So = P = [11] PT = 17 | -1 1].